# QCSE: A Pretrained Quantum Context-Sensitive Word Embedding for Natural Language Processing

### v1

@@ VI Applications and Future Directions
The proposed quantum embedding model enables key NLP applications , such as language modeling, translation, and text generation by capturing linguistic relationships. QCSE can be extended to quantum encoder-decoder frameworks [66, 67, 68, 69] for tasks like translation and summarization. Quantum states encode context-sensitive embeddings, while decoders reconstruct classical outputs, leveraging quantum parallelism for efficiency. Quantum self-attention [68, 69] could compute superposition-based attention scores, reducing the quadratic complexity of classical transformers [63]. Integrating quantum embeddings into pretrained models (e.g., BERT, GPT [70]) may enhance context sensitivity.
Beyond bag-of-words, Skip-Gram [60] or dependency-based contexts could enrich training. Attention mechanisms [63] may dynamically weight context words, enabling quantum transformers with masked/causal language modeling. QCSE benefits low-resource languages [71, 72] (e.g., Fulani) by extracting richer semantics from limited data.
@@
  • Different Ansatz Exploration: Different types of parameterized quantum circuits [73] should be explored. A more suitable ansatz could potentially improve the encoding of intricate word relationships, especially in low-resource settings.
@@ VII Conclusion
This work presented a pretrained Quantum Context-Sensitive Embedding (QCSE) model that leverages quantum computing principles to address NLP challenges. The study demonstrated the feasibility of applying quantum circuits to word embedding tasks, showing strong potential for capturing semantic relationships that classical models may struggle with. The approach encodes word–context relationships as quantum states, exploiting quantum properties such as superposition and entanglement to represent linguistic patterns. Evaluations on both medium- and small-sized datasets show that the model generalizes well, with notable success in low-resource language scenarios. This suggests that quantum models can mitigate the data scarcity problem often faced in NLP, learning meaningful semantic relationships even from limited data. Overall, this work establishes a foundation for further exploration of quantum word embeddings, particularly for low-resource NLP applications.

### v2

@@ VI Applications and Future Directions
The proposed quantum embedding model enables key NLP applications , such as language modeling, translation, and text generation by capturing linguistic relationships. QCSE can be extended to quantum encoder-decoder frameworks [chen2020quantum, bausch2021quantum, zhao2024qksan, shi2024qsan] for tasks like translation and summarization. Quantum states encode context-sensitive embeddings, while decoders reconstruct classical outputs, leveraging quantum parallelism for efficiency. Quantum self-attention [zhao2024qksan, shi2024qsan] could compute superposition-based attention scores, reducing the quadratic complexity of classical transformers [vaswani2017attention]. Integrating quantum embeddings into pretrained models (e.g., BERT, GPT [devlin2018bert]) may enhance context sensitivity. Using the right number of trainable parameters, QCSE could benefit low-resource languages [magueresse2020low, joshi2019unsung] (e.g., Fulani) by extracting richer semantics from limited data.
@@
  • Different Ansatz Exploration: Different types of parameterized quantum circuits [sim2019expressibility] should be explored. A more suitable ansatz could potentially improve the encoding of intricate word relationships, especially in low-resource settings.
@@ VII Conclusion
This work presented a pretrained Quantum Context-Sensitive Embedding (QCSE) model that leverages quantum computing principles to address NLP challenges. The study demonstrated the feasibility of applying quantum circuits to word embedding tasks, showing potential for capturing semantic relationships. The approach encodes word–context relationships as quantum states, exploiting quantum properties such as superposition and entanglement to represent linguistic patterns. Our results show that the proposed QCSE model can capture contextual relationships of texts, even if the provided corpus is limited. Overall, this work contributes to the foundation for further exploration of quantum word embeddings, particularly for low-resource NLP applications.
@@ Acknowledgements
Appendix A Alternative Context Encoding Methods And Results
A-A Alternative Context Encoding Methods
Section IV detailed the exponential decay sinusoidal encoding method that demonstrated optimal performance in experiments among all encoding strategies we tested. In this section, we discuss the results of the alternative strategies. Each method offers distinct characteristics in how it encodes word identity, positional information, and inter-word relationships. This appendix presents these alternative approaches, their motivations, and comparative experimental results.
All encoding methods share the common framework of transforming a context window of $n$ words into a matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ (or vector subsequently reshaped to matrix form), which is then reshaped to $\mathbf{\tilde{C}}\in\mathbb{R}^{2m\times\lceil n^{2}/2m\rceil}$ and padded with zeros as needed for circuit encoding. The fundamental vocabulary-based angle $\theta_{i}=\text{idx}_{i}\cdot\frac{2\pi}{|V|}$ appears across all methods, providing a consistent basis for word identity encoding.
A-A1 Index-Based Diagonal Method
This method introduces asymmetric treatment of diagonal versus off-diagonal elements to emphasize self-relationships while maintaining decay-based encoding for word interactions:
The logarithmic scaling of diagonal elements $\log(1+\text{idx}_{i})$ provides stronger representation for words with higher vocabulary indices. This design assumes that rarer words (typically assigned higher indices in sorted vocabularies) require more prominent self-representation to prevent being overshadowed by more common words. The off-diagonal elements retain exponential distance decay but use only a single sinusoidal term rather than the product $\sin(\omega\theta_{i})\cos(\omega\theta_{j})$ employed in the exponential decay method.
A-A2 Positional Phase Shift Method
This approach emphasizes absolute position encoding alongside word identity through position-dependent phase modulation:
$c_{ij}=e^{-\alpha|i-j|}\sin(\omega i+\delta\theta_{i})+\theta_{i}$ (26)
where $\delta>0$ controls the phase shift scaling. The term $\omega i$ introduces a position-dependent base phase that varies linearly with position in the context window, while $\delta\theta_{i}$ adds word-specific phase modulation. This combination enables the model to distinguish identical words appearing at different positions—a capability that may prove valuable for tasks requiring fine-grained word order sensitivity beyond simple proximity relationships.
A-A3 Hash-Based Modulation Method
To enhance representational uniqueness and reduce potential collisions in the encoding space, this method incorporates a hash function:
$c_{ij}=e^{-\alpha|i-j|}\sin(\omega i+h_{i})+\theta_{i}$ (27)
where $h_{i}=(\text{idx}_{i}\cdot p)\bmod N$ with $p$ being a prime number and $N$ the hash space size. The hash function disperses similar vocabulary indices across the encoding space, potentially improving discrimination when word indices are clustered. The use of prime numbers helps ensure good distributional properties, reducing the likelihood that structurally similar words receive near-identical encodings.
A-A4 Positional Angular Shift Vector Method
Unlike the previous matrix-based methods, this approach begins with a vector representation:
$v_{i}=\omega\theta_{i}$ (28)
producing a context vector $\mathbf{v}\in\mathbb{R}^{n}$ where each element directly encodes the scaled vocabulary angle. This vector is subsequently reshaped into a square matrix for circuit encoding. The method represents the minimal encoding strategy, emphasizing word identity through angular transformation while implicitly encoding position through element ordering. Its simplicity offers computational efficiency but sacrifices the explicit pairwise relationship encoding present in matrix-based methods.
A-B Results
Tables III and IV, and Figures  5,  6 and  7 present the accuracy results and the training dynamics for all encoding methods across different model depths on both datasets.
TABLE III: Accuracy results for alternative context encoding methods on English dataset
Model Params Index Diagonal Phase Shift Hash Mod. Angular Vector
QCSE-1L 21 48.7 54.0 54.0 49.6
QCSE-2L 42 63.7 63.7 62.9 69.1
QCSE-3L 63 66.4 61.1 66.4 65.5
QCSE-4L 84 69.1 71.7 68.1 69.9
QCSE-5L 105 67.3 65.5 65.5 70.8
QCSE-6L 126 67.3 68.1 63.7 69.1
QCSE-7L 147 67.3 69.0 69.1 70.8
QCSE-8L 168 80.5 85.0 81.4 69.1
TABLE IV: Accuracy results for alternative context encoding methods on Fulani dataset
Model Params Index Diagonal Phase Shift Hash Mod. Angular Vector
QCSE-1L 15 38.5 50.0 42.3 42.3
QCSE-2L 30 38.5 30.8 53.8 53.8
QCSE-3L 45 53.8 53.8 46.2 53.8
QCSE-4L 60 65.4 69.2 61.5 50.0
QCSE-5L 75 65.4 65.4 73.1 57.7
QCSE-6L 90 61.5 69.4 65.4 57.7
QCSE-7L 105 76.9 84.6 50.0 61.5
QCSE-8L 120 65.4 76.9 73.1 57.7
(a) Trainable parameter count vs. Accuracy using English dataset
(b) Trainable parameter count vs. Accuracy using Fulani dataset
Figure 5: Accuracy versus trainable parameter count in QCSE. The number of trainable parameters influences performance.
(a) QCSE with 1 layer
(b) QCSE with 2 layers
(c) QCSE with 3 layers
(d) QCSE with 4 layers
(e) QCSE with 5 layers
(f) QCSE with 6 layers
(g) QCSE with 7 layers
(h) QCSE with 8 layers
Figure 6: Training loss progression for QCSE models with exponential decay sinusoidal encoding across different depths (1-8 layers) on the English dataset.
(a) QCSE with 1 layer
(b) QCSE with 2 layers
(c) QCSE with 3 layers
(d) QCSE with 4 layers
(e) QCSE with 5 layers
(f) QCSE with 6 layers
(g) QCSE with 7 layers
(h) QCSE with 8 layers
Figure 7: Training loss plots of QCSE using same number of layers with different context-encoding methods on Fulani dataset.
The alternative encoding methods exhibit variable and comparable performance compared to the exponential decay sinusoidal approach. On the English dataset, peak accuracies range from 69.1-85.0% (compared to 85.8% for exponential decay), with optimal depths varying by method: Index-Based Diagonal Method peaks at 8 layers (80.5%), Positional Phase Shift Method at 8 layers (85.0%), Hash-Based Modulation Method at 8 layers (81.4%), and Positional Angular Shift Vector Method plateaus at 5-7 layers (70.8%). This variability means that the encoding strategy influences the performance of the quantum circuit and that larger and more diverse datasets are necessary for reliably discriminating between encoding strategies in quantum embedding models.
The Positional Angular Shift Vector Method demonstrates limited performance, maintaining 69-71% accuracy across 2-8 layers on English data. This reflects its simpler encoding structure—without explicit pairwise relationship terms, the model reaches its representational limit early and cannot benefit from additional circuit depth. In contrast, Index Diagonal and Hash Modulation methods show continued improvement through 8 layers, suggesting that their encodings provide sufficient representational to exploit deeper circuits, though outperformed by the exponential decay method.
On the smaller Fulani dataset, some alternative methods exhibit instability. The Hash-Based Modulation method ranges from 42.3% (1L) to 73.1% (5L and 8L), with a collapse at 7 layers (50.0%). The Positional Phase Shift method achieves the highest peak (84.6% at 7L) but shows variation across depths (30.8% at 2L). This instability contrasts with the exponential decay method progression, suggesting that the product term $\sin(\omega\theta_{i})\cos(\omega\theta_{j})$ in Equation 24 provides robust optimization properties. The Index-Based Diagonal encoding using a self-relationships (via $\log(1+\text{idx}_{i})$ diagonal terms) achieves comparable performance (80.5% at 8L on English) but does not outperform the exponential decay approach. The Positional Phase Shift method employing an emphasis on absolute position (through $\omega i$) yields the second-best overall result (85.0%), nearly matching exponential decay, suggesting that position encoding complementary to relative distance may merit further investigation. The Hash-Based Modulation mechanism shows promise (81.4%) but introduces additional hyperparameter sensitivity (prime $p$, hash space $N$) not present in exponential decay.
Figures 6 and 7 show depth-dependent training dynamics across encoding methods. Initial loss increases with circuit depth (5.0 for 1 layer to 6.2–6.9 for 7–8 layers), reflecting the higher-dimensional optimization landscape of deeper circuits. Three convergence regimes emerge: shallow models (1–2 layers) exhibit minimal loss reduction and early plateau, leading to poor accuracy; medium-depth models (3–6 layers) show steady descent and moderate performance; and deep models (7–8 layers) display more loss reduction, achieving the highest accuracies. Overall, effective learning in quantum embedding models is shown across all the context encoding methods.